TRUE/FALSE: A line with a slope 1 alaways passes through the origin.
If false, please give a counterexample. Thank you!

Answer by swincher4391(1107)

FALSE. Given a point (m,n), where m and n are not equal, and a slope 1, we can come up with an equation using the point-slope form.
Let's choose the point (1,2).
Then y-y1 = m(x-x1)
y-2 = 1(x-1)
y -2 = x-1
y = x + 1 <--- COUNTEREXAMPLE
That is an equation of a line with slope 1 and does not pass through the origin.